# Robust $\Lambda$-quantiles and extreme probabilities

### v1

@@ 4.1 Moment constraints
For an increasing $\Lambda$, it can be observed from Table 1 that $q_{\Lambda}^{-}$ is strictly smaller than $\tilde{q}_{\Lambda}^{-}$, and $q_{\Lambda}^{+}$ is strictly smaller than $\tilde{q}_{\Lambda}^{+}$. This verifies the conclusion in (i) of Proposition 2. Note that the bold numbers in Table 1 may not be the true best-case or worst-case values; instead, they represent the lower bounds of the worst-case values or the upper bounds of the best-case values, respectively. Moreover, we observe that $\tilde{q}_{\Lambda}^{-}$, $\tilde{q}_{\Lambda}^{+}$, and the worst-case $\Lambda$-quantiles in Table 1 are consistently larger than those in Table 2. This shows that $\Lambda$-quantiles with an increasing $\Lambda$ tend to penalize more on scenarios with significant capital losses.
@@
Our results in Theorem 2 can be applied to the robust portfolio selection. Let $\mathbf{X}=(X_{1},...,X_{n})\in\mathcal{X}^{n}$ represent the loss or negative return for $n$ different assets. The set of all possible portfolio positions are denoted by $\Delta_{n}=\left\{\mathbf{w}=\left(w_{1},\ldots,w_{n}\right)\in[0,1]^{n}:\sum_{i=1}^{n}w_{i}=1\right\}$. Note that here short-selling is not allowed. For $\mathbf{w}\in\Delta_{n}$, the risk of the portfolio $\mathbf{w}^{\top}\mathbf{X}$ is evaluated by $\rho\left(F_{\mathbf{w}^{\top}\mathbf{X}}\right)$, where $\rho$ is some risk measure. Suppose we know the mean and the upper bound of the covariance matrix of $\mathbf{X}$. Then the uncertainty set of the portfolio is given by
@@
$\mathcal{M}_{2}(\mathbf{w},\boldsymbol{\mu},\Sigma)=\left\{F_{\mathbf{w}^{\top}\mathbf{X}}\in\mathcal{M}_{2}:\mathbb{E}[\mathbf{X}]=\boldsymbol{\mu},\operatorname{Cov}(\mathbf{X})\preceq\Sigma\right\},$ (6)
@@
where $\Sigma$ is a semidefinite symmetric matrix and for a semidefinite symmetric matrix $B$, $B\preceq\Sigma$ means $\Sigma-B$ is positive semidefinite. It follows from the result of Popescu (2007) that $\mathcal{M}_{2}(\mathbf{w},\boldsymbol{\mu},\Sigma)=\mathcal{M}_{2}(\mathbf{w}^{\top}\boldsymbol{\mu},\sqrt{\mathbf{w}^{\top}\Sigma\mathbf{w}})$. The optimal portfolio selection with known mean and upper bound of covariance matrix is formulated as below
@@
$\min_{\mathbf{w}\in{\Delta_{n}}}\sup_{F\in\mathcal{M}_{2}(\mathbf{w},\boldsymbol{\mu},\Sigma)}\rho\left(F\right).$
@@
We refer to Pflug and Pohl (2018) for an overview of the portfolio selection with model uncertainty. Note that return of the portfolio can also be considered as a constraint in the above optimization problem. For instance, we can impose the constraint $\sum_{i=1}^{n}w_{i}\mu_{i}\leqslant c$ for some $c<0$ on the above optimization problem, requiring that the expected return of the portfolio is larger than $-c$. Then the optimization problem becomes $\min_{\mathbf{w}\in{\Delta_{n}^{\prime}}}\sup_{F\in\mathcal{M}_{2}(\mathbf{w},\boldsymbol{\mu},\Sigma)}\rho\left(F\right)$ with $\Delta_{n}^{\prime}=\left\{\mathbf{w}\in\Delta_{n}:\sum_{i=1}^{n}w_{i}\mu_{i}\leqslant c\right\}$, which does not change the nature of the problem. To simplify the problem, we here consider the portfolio selection problem with $c\geqslant\max_{i=1}^{n}\mu_{i}$.
@@
Let $\mathcal{M}_{2}(\mathbf{w},\boldsymbol{\mu},\Sigma)$ be given in (6) and $\Lambda:\mathbb{R}\to[0,1]$. For $\rho=\tilde{q}_{\Lambda}^{+}$ with decreasing $\Lambda$ or $\rho=\tilde{q}_{\Lambda}^{-}$, we have
$\min_{\mathbf{w}\in{\Delta_{n}}}\sup_{F\in\mathcal{M}_{2}(\mathbf{w},\boldsymbol{\mu},\Sigma)}\rho\left(F\right)=\min_{\mathbf{w}\in{\Delta_{n}}}\rho\left(F^{-}_{\mathcal{M}_{2}(\mathbf{w}^{\top}\boldsymbol{\mu},\sqrt{\mathbf{w}^{\top}\Sigma\mathbf{w}})}\right),$ (7)
@@
$F^{-}_{\mathcal{M}_{2}(\mathbf{w},\boldsymbol{\mu},\Sigma)}(x)=\frac{(x-\mathbf{w}^{\top}\boldsymbol{\mu})^{2}}{(\mathbf{w}^{\top}\Sigma\mathbf{w}+(x-\mathbf{w}^{\top}\boldsymbol{\mu})^{2})}\mathds{1}_{\{x\geqslant\mathbf{w}^{\top}\boldsymbol{\mu}\}}.$
@@
$\sup_{F\in\mathcal{M}_{2}(\mathbf{w},\boldsymbol{\mu},\Sigma)}\rho\left(F\right)=\rho\left(F^{-}_{\mathcal{M}_{2}(\mathbf{w}^{\top}\boldsymbol{\mu},\sqrt{\mathbf{w}^{\top}\Sigma\mathbf{w}})}\right),$
@@
$\min_{\mathbf{w}\in{\Delta_{n}}}\sup_{F\in\mathcal{M}_{2}(\mathbf{w},\boldsymbol{\mu},\Sigma)}\rho\left(F\right)=\min_{\mathbf{w}\in{\Delta_{n}}}\rho\left(F^{-}_{\mathcal{M}_{2}(\mathbf{w}^{\top}\boldsymbol{\mu},\sqrt{\mathbf{w}^{\top}\Sigma\mathbf{w}})}\right).$
@@
Here, we consider only the worst case of $\tilde{q}^{-}_{\Lambda}$, as $\sup_{F\in\mathcal{M}_{p}(m,v)}\tilde{q}_{\Lambda}^{-}(F)=\tilde{q}_{\Lambda}^{-}(F_{\mathcal{M}_{p}(m,v)}^{-})$ holds for both increasing and decreasing $\Lambda$.
Figures 2 and 2 demonstrate the worst-case values of the portfolio as $w_{1}$ varies, showing that the optimal portfolio positions under the $\Lambda$-quantile criterion can differ from those under the VaR criterion. Additionally, for an increasing $\Lambda$ in Figure 2, when the mean of the portfolio is relatively small, the DM uses the small probability level of $0.8$ to determine the portfolio positions; when the mean of the portfolio is relatively large, a larger probability level of $0.95$ is used. But if $\Lambda$ is a decreasing function in Figure 2, this situation is reversed. This phenomenon is due to the fact that an increasing $\Lambda$ may penalize large losses, while a decreasing $\Lambda$ suggests that the DM accepts a smaller probability for larger potential losses, indicating a relative preference for risk. Moreover, when two assets are positively correlated, the worst-case values are higher compared to the case with negatively correlated assets, as negative correlation results in hedging effect.
Figure 1: The worst-case values of $\tilde{q}^{-}_{\Lambda}$ and $q^{-}_{\lambda}$ in $w_{1}$ with $\Lambda(\alpha,\beta;z)=\Lambda(0.8,0.95;3)$ (blue solid line) and $\lambda=0.9$ (black dashed line)
Figure 2: The worst-case values of $\tilde{q}^{-}_{\Lambda}$ and $q^{-}_{\lambda}$ in $w_{1}$ with $\Lambda(\alpha,\beta;z)=\Lambda(0.95,0.8;6)$ (blue solid line) and $\lambda=0.9$ (black dashed line)
@@ 4.2 Probability constraints via Wasserstein distance
where the parameter $\varepsilon$ represents the magnitude of uncertainty. Let $x_{+}=\max(x,0)$. In the following result, we will give the expressions of $F_{\mathcal{M}_{p}(G,\varepsilon)}^{-}$ and $F_{\mathcal{M}_{p}(G,\varepsilon)}^{+}$.
@@
where for $\alpha\in(0,1)$, $l(\alpha)$ is the unique solution of the following equation
@@
By (i) of Corollary 1, we have $\sup_{F\in\mathcal{M}_{p}(G,\varepsilon)}q_{\alpha}^{-}(F)=q_{\alpha}^{-}(F_{\mathcal{M}_{p}(G,\varepsilon)}^{-})=l(\alpha)$. For $10<\alpha_{1}<\alpha_{2}<$ 1, it follows that $\int_{\alpha_{2}}^{1}\left(l(\alpha_{1})-q^{-}_{t}(G)\right)_{+}^{p}\mathrm{~{}d}t<\varepsilon^{p}$. This implies $l(\alpha_{2})>l(\alpha_{1})$. Hence, $l$ is strictly increasing over $(0,1)$, which implies $F_{\mathcal{M}_{p}(G,\varepsilon)}^{-}(x)=l_{+}^{-1}(x)$ for $x\in\mathbb{R}$ and $F_{\mathcal{M}_{p}(G,\varepsilon)}^{-}$ is continuous.
@@
where $\delta_{x}$ is the distribution with probability mass 1 on $x$. By Chebyshev’s inequality, we have $\sup_{k\geqslant 1}(1-F_{k}(x))\leqslant\frac{\left(W_{p}(\delta_{0},G)+\varepsilon\right)^{p}}{x^{p}}$ for $x>0$ and $\sup_{k\geqslant 1}F_{k}(x)\leqslant\frac{\left(W_{p}(\delta_{0},G)+\varepsilon\right)^{p}}{|x|^{p}}$ for $x<0$. Hence, $\{F_{k},~{}k\geqslant 1\}$ is tight. By Theorem 25.10 of Billingsley (1995)) that there exist a subsequence $\{F_{k_{l}},~{}l\geqslant 1\}$ and a distribution $F$ such that $F_{k_{l}}\to F$ weakly as $l\to\infty$, which implies $F_{k_{l}}^{-1}(t)\to F^{-1}(t)$ over $(0,1)$ a.s.. It follows from Fatou’s lemma that
@@
Following Theorem 1, Proposition 3, and Proposition 8, we immediately obtain the following conclusion.
@@
In the following example, we consider the uncertainty set $\mathcal{M}_{p}(G,\varepsilon)$ with $G$ being normal, exponential and uniform distributions. Note that $F_{\mathcal{M}_{p}(m,v)}^{-}$ and $F_{\mathcal{M}_{p}(m,v)}^{+}$ can be computed numerically using (11). For decreasing $\Lambda$, it is more convenient to calculate the best-case and worst case of $\Lambda$quantiles using the expressions in Proposition 5 as we do not need to compute the inverse functions of $l$ and $u$.
@@
Let $p=1$, $\varepsilon=0.1$ and $\Lambda_{(\alpha,\beta;z)}:x\mapsto\alpha\mathds{1}_{\{x<z\}}+\beta\mathds{1}_{\{x\geqslant z\}}$. Applying the results in Theorem 3 and Proposition 5, we obtain the robust $\Lambda$-quantiles numerically, which are displayed in Tables 3 and 4. Again, the bold numbers in Table 3 represent the upper bounds for the best values and the lower bounds for the worst values.
We observe that the four $\Lambda$-quantiles in this example are all the same. Also, similar to Example 4, worst-case values of $\tilde{q}_{\Lambda}^{-}$ with an increasing $\Lambda$ are larger than those with a decreasing $\Lambda$. This indicates that $\tilde{q}_{\Lambda}^{-}$ with increasing $\Lambda$ functions are more conservative under this uncertainty set.
@@
Based on Lemma 1, the optimization problems (12) can be solved using Theorem 3.
@@
Figures 4 and 4 show the worst-case values of the portfolio varying in $w_{1}$. The optimal portfolio positions under the $\Lambda\mathrm{VaR}$ criterion can be different from those under the VaR criterion. In Figure 4, with an increasing $\Lambda$, when the mean of the benchmark distribution is relatively small, the DM uses the small probability level of $0.8$ to determine the portfolio positions; when the mean of the portfolio is relatively large, a larger probability level of $0.95$ is used. But in Figure 4, where $\Lambda$ is a decreasing function, this situation is reversed. Additionally, the optimal portfolio positions are higher when the covariance matrix of the benchmark distribution is positively correlated compared to the case with negatively correlated assets. All these phenomena can be explained similarly to Example 5.
Figure 3: The worst-case values of $\tilde{q}^{-}_{\Lambda}$ and $q^{-}_{\lambda}$ in $w_{1}$ with $\Lambda(\alpha,\beta;z)=\Lambda(0.8,0.95;6)$ (blue solid line) and $\lambda=0.9$ (black dashed line)
Figure 4: The worst-case values of $\tilde{q}^{-}_{\Lambda}$ and $q^{-}_{\lambda}$ in $w_{1}$ with $\Lambda(\alpha,\beta;z)=\Lambda(0.95,0.8;6)$ (blue solid line) and $\lambda=0.9$ (black dashed line)

### v2

@@ 4.1 Moment constraints
For an increasing $\Lambda$, it can be observed from Table 1 that $q_{\Lambda}^{-}$ is strictly smaller than $\tilde{q}_{\Lambda}^{-}$, and $q_{\Lambda}^{+}$ is strictly smaller than $\tilde{q}_{\Lambda}^{+}$. This finding verifies the conclusion in (i) of Proposition 2. Note that the bold numbers in Table 1 may not be the true best-case or worst-case values; instead, they represent the lower bounds of the worst-case values or the upper bounds of the best-case values, respectively. Moreover, we observe that $\tilde{q}_{\Lambda}^{-}$, $\tilde{q}_{\Lambda}^{+}$, and the worst-case $\Lambda$-quantiles in Table 1 are consistently larger than those in Table 2. This shows that $\Lambda$-quantiles with an increasing $\Lambda$ tend to penalize more in scenarios with significant capital losses. Finally, we observe from Table 2 that the numerical values of the four $\Lambda$-quantiles are the same. In fact, their theoretical values are also the same. This is because of the special choice of $\Lambda$ and distribution functions. In general, those four values are not the same; see e.g., Examples 4 and 5 in Bellini and Peri (2022).
@@
Our results in Theorem 2 can be applied to robust portfolio selection. Let $\mathbf{X}=(X_{1},...,X_{n})\in\mathcal{X}^{n}$ represent the loss or negative return for $n$ different assets. The set of all possible portfolio positions is denoted by $\Delta_{n}=\left\{\mathbf{w}=\left(w_{1},\ldots,w_{n}\right)\in[0,1]^{n}:\sum_{i=1}^{n}w_{i}=1\right\}$. Note that here short-selling is not allowed. For $\mathbf{w}\in\Delta_{n}$, the risk of the portfolio $\mathbf{w}^{\top}\mathbf{X}$ is evaluated by $\rho\left(F_{\mathbf{w}^{\top}\mathbf{X}}\right)$, where $\rho$ is some risk measure. Suppose that we know the mean and the upper bound of the covariance matrix of $\mathbf{X}$. Then the uncertainty set of the portfolio is given by
@@
^$\widehat{\mathcal{M}}(\mathbf{w},\boldsymbol{\mu},\Sigma)=\left\{F_{\mathbf{w}^{\top}\mathbf{X}}\in\mathcal{M}_{2}:\mathbb{E}[\mathbf{X}]=\boldsymbol{\mu},\operatorname{Cov}(\mathbf{X})\preceq\Sigma\right\},$ (6)
@@
where $\Sigma$ is a semidefinite symmetric matrix and for a semidefinite symmetric matrix $B$, $B\preceq\Sigma$ means $\Sigma-B$ is positive semidefinite. It follows from the result of Popescu (2007) that ^$\widehat{\mathcal{M}}(\mathbf{w},\boldsymbol{\mu},\Sigma)=\mathcal{M}_{2}(\mathbf{w}^{\top}\boldsymbol{\mu},\sqrt{\mathbf{w}^{\top}\Sigma\mathbf{w}})$. The optimal portfolio selection with known mean and upper bound of the covariance matrix is formulated as follows
@@
$\min_{\mathbf{w}\in{\Delta_{n}}}\sup_{F\in\widehat{\mathcal{M}}(\mathbf{w},\boldsymbol{\mu},\Sigma)}\rho\left(F\right).$
@@
We refer to Pflug and Pohl (2018) for an overview of the portfolio selection with model uncertainty. Note that return of the portfolio can also be considered as a constraint in the above optimization problem. For instance, we can impose the constraint $\sum_{i=1}^{n}w_{i}\mu_{i}\leqslant c$ for some $c<0$ on the above optimization problem, requiring that the expected return of the portfolio is larger than $-c$. Then the optimization problem becomes $\min_{\mathbf{w}\in{\Delta_{n}^{\prime}}}\sup_{F\in\widehat{\mathcal{M}}(\mathbf{w},\boldsymbol{\mu},\Sigma)}\rho\left(F\right)$ with $\Delta_{n}^{\prime}=\left\{\mathbf{w}\in\Delta_{n}:\sum_{i=1}^{n}w_{i}\mu_{i}\leqslant c\right\}$, which does not change the nature of the problem. To simplify the problem, we here consider the portfolio selection problem with $c\geqslant\max_{i=1}^{n}\mu_{i}$.
@@
Let ^$\widehat{\mathcal{M}}(\mathbf{w},\boldsymbol{\mu},\Sigma)$ be given in (6) and $\Lambda:\mathbb{R}\to[0,1]$. For $\rho=\tilde{q}_{\Lambda}^{+}$ with decreasing $\Lambda$ or $\rho=\tilde{q}_{\Lambda}^{-}$, we have
$\min_{\mathbf{w}\in{\Delta_{n}}}\sup_{F\in\widehat{\mathcal{M}}(\mathbf{w},\boldsymbol{\mu},\Sigma)}\rho\left(F\right)=\min_{\mathbf{w}\in{\Delta_{n}}}\rho\left(F^{-}_{\mathcal{M}_{2}(\mathbf{w}^{\top}\boldsymbol{\mu},\sqrt{\mathbf{w}^{\top}\Sigma\mathbf{w}})}\right),$ (7)
@@
$F^{-}_{\widehat{\mathcal{M}}(\mathbf{w},\boldsymbol{\mu},\Sigma)}(x)=\frac{(x-\mathbf{w}^{\top}\boldsymbol{\mu})^{2}}{(\mathbf{w}^{\top}\Sigma\mathbf{w}+(x-\mathbf{w}^{\top}\boldsymbol{\mu})^{2})}\mathds{1}_{\{x\geqslant\mathbf{w}^{\top}\boldsymbol{\mu}\}}.$
@@
$\sup_{F\in\widehat{\mathcal{M}}(\mathbf{w},\boldsymbol{\mu},\Sigma)}\rho\left(F\right)=\rho\left(F^{-}_{\mathcal{M}_{2}(\mathbf{w}^{\top}\boldsymbol{\mu},\sqrt{\mathbf{w}^{\top}\Sigma\mathbf{w}})}\right),$
@@
$\min_{\mathbf{w}\in{\Delta_{n}}}\sup_{F\in\widehat{\mathcal{M}}(\mathbf{w},\boldsymbol{\mu},\Sigma)}\rho\left(F\right)=\min_{\mathbf{w}\in{\Delta_{n}}}\rho\left(F^{-}_{\mathcal{M}_{2}(\mathbf{w}^{\top}\boldsymbol{\mu},\sqrt{\mathbf{w}^{\top}\Sigma\mathbf{w}})}\right).$
@@
Here, we consider the worst-case values of $\tilde{q}^{-}_{\Lambda}$, as $\sup_{F\in\mathcal{M}_{p}(m,v)}\tilde{q}_{\Lambda}^{-}(F)=\tilde{q}_{\Lambda}^{-}(F_{\mathcal{M}_{p}(m,v)}^{-})$ holds for both increasing and decreasing $\Lambda$. As a comparison, we also consider the worst-case values of $q_{\alpha}^{-}$ for some $\alpha\in(0,1)$
Figures 2 and 2 demonstrate the worst-case values of the portfolio as functions of $w_{1}$, showing that the optimal portfolio positions under the $\Lambda$-quantile criterion can differ from those under the VaR criterion. Additionally, for an increasing $\Lambda$ in Figure 2, when the mean of the portfolio is relatively small, the DM uses a small probability level of $0.8$ to determine the portfolio positions; when the mean of the portfolio is relatively large, a larger probability level of $0.95$ is used. However, if $\Lambda$ is a decreasing function in Figure 2, this situation is reversed. This phenomenon is due to the fact that an increasing $\Lambda$ may penalize large losses, while a decreasing $\Lambda$ suggests that the DM accepts a higher probability for larger potential losses. Moreover, when two assets are positively correlated, the worst-case values are higher compared to the case with negatively correlated assets, as a negative correlation results in a hedging effect.
Figure 1: The worst-case values of $\tilde{q}^{-}_{\Lambda}$ with $\Lambda=\Lambda_{(0.8,0.95;3)}$ (blue solid line) and $q^{-}_{\alpha}$ with $\alpha=0.9$ (black dashed line) as functions of $w_{1}$
Figure 2: The worst-case values of $\tilde{q}^{-}_{\Lambda}$ with $\Lambda=\Lambda_{(0.95,0.8;6)}$ (blue solid line) and $q^{-}_{\alpha}$ with $\alpha=0.9$ (black dashed line) as functions of $w_{1}$
@@ 4.2 Probability constraints via Wasserstein distance
where the parameter $\varepsilon$ represents the magnitude of uncertainty. In the following result, we will give the expressions of $F_{\mathcal{M}_{p}(G,\varepsilon)}^{-}$ and $F_{\mathcal{M}_{p}(G,\varepsilon)}^{+}$.
@@
where for each $\alpha\in(0,1)$, $l(\alpha)$ is the unique solution of the following equation
@@
By (i) of Corollary 1, we have $\sup_{F\in\mathcal{M}_{p}(G,\varepsilon)}q_{\alpha}^{-}(F)=q_{\alpha}^{-}(F_{\mathcal{M}_{p}(G,\varepsilon)}^{-})=l(\alpha)$. For $10<\alpha_{1}<\alpha_{2}<$ 1, it follows that $\int_{\alpha_{2}}^{1}\left(l(\alpha_{1})-q^{-}_{t}(G)\right)_{+}^{p}\mathrm{~{}d}t<\varepsilon^{p}$. This implies that $l(\alpha_{2})>l(\alpha_{1})$. Hence, $l$ is strictly increasing over $(0,1)$, which implies that $F_{\mathcal{M}_{p}(G,\varepsilon)}^{-}(x)=l_{+}^{-1}(x)$ for $x\in\mathbb{R}$ and $F_{\mathcal{M}_{p}(G,\varepsilon)}^{-}$ is continuous.
@@
where $\delta_{x}$ is the distribution with probability mass 1 on $x$. By Chebyshev’s inequality, we have $\sup_{k\geqslant 1}(1-F_{k}(x))\leqslant\frac{\left(W_{p}(\delta_{0},G)+\varepsilon\right)^{p}}{x^{p}}$ for $x>0$ and $\sup_{k\geqslant 1}F_{k}(x)\leqslant\frac{\left(W_{p}(\delta_{0},G)+\varepsilon\right)^{p}}{|x|^{p}}$ for $x<0$. Hence, $\{F_{k},~{}k\geqslant 1\}$ is tight. By Theorem 25.10 of Billingsley (1995)), there exist a subsequence $\{F_{k_{l}},~{}l\geqslant 1\}$ and a distribution $F$ such that $F_{k_{l}}\to F$ weakly as $l\to\infty$, which implies $F_{k_{l}}^{-1}(t)\to F^{-1}(t)$ over $(0,1)$ a.s.. It follows from Fatou’s lemma that
@@
Following Theorem 1, Proposition 3, and Proposition 8, we immediately obtain the following conclusions.
@@
In the following example, we consider the uncertainty set $\mathcal{M}_{p}(G,\varepsilon)$ with $G$ being normal, exponential and uniform distributions. Note that $F_{\mathcal{M}_{p}(G,\varepsilon)}^{-}$ and $F_{\mathcal{M}_{p}(G,\varepsilon)}^{+}$ can be computed numerically via (11). For decreasing $\Lambda$, it is more convenient to calculate the best-case and worst case of $\Lambda$quantiles using the expressions in Proposition 5 as we do not need to compute the inverse functions of $l$ and $u$.
@@
Let $p=1$, $\varepsilon=0.1$ and $\Lambda_{(\alpha,\beta;z)}:x\mapsto\alpha\mathds{1}_{\{x<z\}}+\beta\mathds{1}_{\{x\geqslant z\}}$. Applying the results in Theorem 3 and Proposition 5, we obtain the robust $\Lambda$-quantiles numerically, which are displayed in Tables 3 and 4. Again, the bold numbers in Table 3 represent the upper bounds for the worst-case values and the lower bounds for the best-case values.
We observe that the values and bounds of the four $\Lambda$-quantiles displayed in Tables 3 and 4 are all the same. Moreover, the worst-case values of $\tilde{q}_{\Lambda}^{-}$ with an increasing $\Lambda$ are larger than those with a decreasing $\Lambda$. This finding indicates that $\tilde{q}_{\Lambda}^{-}$ with increasing $\Lambda$ functions are more conservative under this uncertainty set.
@@
Based on Lemma 1, the optimization problem (12) can be solved using Theorem 3.
@@
Figures 4 and 4 show the worst-case values of the portfolio as functions of $w_{1}$. The optimal portfolio positions under the $\Lambda$-quantile criterion can differ from those under the VaR criterion. In Figure 4, with an increasing $\Lambda$, when the mean of the benchmark distribution is relatively small, the DM uses the small probability level of $0.8$ to determine the portfolio positions; when the mean of the portfolio is relatively large, a larger probability level of $0.95$ is used. But in Figure 4, where $\Lambda$ is a decreasing function, this situation is reversed. Additionally, the optimal portfolio positions are higher when the covariance matrix of the benchmark distribution is positively correlated compared to the case with negatively correlated assets. All these phenomena can be explained similarly to Example 5.
Figure 3: The worst-case values of $\tilde{q}^{-}_{\Lambda}$ with $\Lambda=\Lambda_{(0.8,0.95;6)}$ (blue solid line) and $q^{-}_{\alpha}$ with $\alpha=0.9$ (black dashed line) as functions of $w_{1}$
Figure 4: The worst-case values of $\tilde{q}^{-}_{\Lambda}$ with $\Lambda=\Lambda_{(0.95,0.8;6)}$ (blue solid line) and $q^{-}_{\alpha}$ with $\alpha=0.9$ (black dashed line) as functions of $w_{1}$